OpenStudy (anonymous):

A particle is moving along the x-axis so that its position at t ≥ 0 is given by s(t) = (t)In(3t). Find the acceleration of the particle when the velocity is first zero

jimthompson5910 (jim_thompson5910):

are you able to take the derivative?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

what do you get when you take the derivative

OpenStudy (anonymous):

ln(t)+ln(3)+1

jimthompson5910 (jim_thompson5910):

which is the same as ln(3t) + 1, good

jimthompson5910 (jim_thompson5910):

set s ' (t) equal to zero and solve for t s ' (t) = 0 ln(3t) + 1 = 0 t = ??

OpenStudy (anonymous):

t=-.9102

jimthompson5910 (jim_thompson5910):

I'm getting a positive value

jimthompson5910 (jim_thompson5910):

Try again

OpenStudy (anonymous):

I got the same thing. I subtracted the -1 to the other side then divided by ln3 that gave me what i have

jimthompson5910 (jim_thompson5910):

ln(3t) + 1 = 0 ln(3t) = -1 3t = e^(-1) t = ??

OpenStudy (anonymous):

.1226

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

this is when the velocity is zero

jimthompson5910 (jim_thompson5910):

now you need to find s '' (t), the acceleration function

OpenStudy (anonymous):

how do i do that?

OpenStudy (anonymous):

do i take the derivative of the derivative?

jimthompson5910 (jim_thompson5910):

yes, you would take the derivative of s ' (t)

OpenStudy (anonymous):

would it be ln3/x?

jimthompson5910 (jim_thompson5910):

I'm getting s '' (t) = 1/t

jimthompson5910 (jim_thompson5910):

be sure to simplify fully (and use the chain rule)

OpenStudy (anonymous):

im confused. what did you do exactly?

jimthompson5910 (jim_thompson5910):

s ' (t) = ln(3t) + 1 s '' (t) = 1/(3t)*3 s '' (t) = 1/t

jimthompson5910 (jim_thompson5910):

I derived ln(3t) to get 1/(3t) and then you have to multiply by the derivative of the inside stuff (the derivative of 3t, which is 3) so using the chain rule if y = ln(3t), then dy/dt = 1/(3t)*3 = 1/t

OpenStudy (anonymous):

Okay so then you have 1/t do I just fill in what i got for t earlier into that formula?

jimthompson5910 (jim_thompson5910):

"Find the acceleration of the particle when the velocity is first zero"

jimthompson5910 (jim_thompson5910):

so you plug t = 0.1226 into s '' (t) to get the acceleration at t = 0.1226

OpenStudy (anonymous):

so 8.1548?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

thats not it is it?

OpenStudy (anonymous):

it gave me answer choices and they are as followed a. 3e^2 b. e c. 3e d. none of the above im guessing that means its d

jimthompson5910 (jim_thompson5910):

hmm one sec

jimthompson5910 (jim_thompson5910):

let's go back to ln(3t) + 1 = 0

jimthompson5910 (jim_thompson5910):

ln(3t) + 1 = 0 ln(3t) = -1 3t = e^(-1) 3t = 1/e t = 1/(3e) so at exactly t = 1/(3e), the particle has velocity 0

jimthompson5910 (jim_thompson5910):

now plug t = 1/(3e) into the acceleration function s '' (t) s '' (t) = 1/t s '' (1/(3e)) = 1/[ 1/(3e) ] s '' (1/(3e)) = 3e so the acceleration is exactly 3e

OpenStudy (anonymous):

Okay then that makes more sense thank you!

jimthompson5910 (jim_thompson5910):

you're welcome